Philosophy 29: Critical Reasoning

Exam, Form: A Name:

Student Number:

TA:

Date:

Section 1. Fill in the blank (5 pts each)

Here is a plaçe where you can put some instructions, so that the students won’t get confused when they see a piece of paper with a whole bunch of questions on it. Of course, the instructions aren’t required to be that lengthy, but you can make them as lengthy as you want so that people know what you are talking about.

1. One main component of Nietzche’s moral philosophy is the .

2. thought that without a strong, centralized, effective government, chaos would reign in the state of nature.

3.  ’s first work was the Tractatus-  Philosophicus.

4. Mill’s theory of morality is known as

5. According to Kant, we should always always always follow the imperative.

6. How much  would a  chuck, if a  would  , wood?

Section 2. True/False (5 pts each)

Ho, hum, another batch of instructions. But this time we will make them a little shorter, because we assume that students realize what they should do.

All animals are created equal, but some animals are more equal than others.

Laden swallows fly faster than unladen swallows, unless they carry coconuts.

‘Monty Python and the Holy Grail’ is a very funny movie.

This sentence is not false.

‘Roger & Me’ chronicles one man’s attempt to get into Disneyland so that he can visit Toontown.

Section 3.

These are meant to be multiple-choice questions — the type you would give students to fill out using scantrons (or whatever you call the number-2 pencil automatic grading machines).

1.

How many people live in Wales?

(a) Approximately 2,811,865.

(b) More than in most countries.

(c) None.

(d) Exactly seventeen.

2.

What philosopher declared that human beings are the conduit through which language passes?

(a) Martha Stewart

(b) Jacques Derrida

(c) Jacques Cousteau

(d) Martin Heidegger 3.

The official state motto of Idaho is?

(a) “We’re friendly than other states.”

(b) “Home of the Unabomber.”

(c) “Famous Potatoes.”

(d) “Try us, you’ll like us.” 4.

How many cows does it take to graze a field?

(a) One.

(b) Two.

(c) Three.

(d) Four

Section 4. Short Answer (10 pts each)

This is an example of the shortanswer type of question, where the questions are rearranged between tests.

1. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back.

2. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy.

3. ‘Beware the Jabberwock, my son! The jaws that bite, the claws that catch! Beware the Jubjub bird, and shun the frumious Bandersnatch!’

4. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; All mimsy were the borogoves, and the mome raths outgrabe.

5. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe.

6. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came!

7. He took his vorpal sword in hand: long time the manxome foe he sought – so rested he by the Tumtum tree, and stood awhile in thought.

Section 5. Essay Questions (50 pts each)

This is an example of the fixed type of question, where the questions aren’t rearranged between tests. Perfect for essay questions, which you normally put at the end and don’t need to rearrange.

1. Question 1.

2. Question 2.

3. Question 3.

4. Question 4.

5. Question 5.

Answer Key for Exam A

Section 1. Fill in the blank (5 pts each)

Here is a plaçe where you can put some instructions, so that the students won’t get confused when they see a piece of paper with a whole bunch of questions on it. Of course, the instructions aren’t required to be that lengthy, but you can make them as lengthy as you want so that people know what you are talking about.

1. One main component of Nietzche’s moral philosophy is the will to power.

2. Hobbes thought that without a strong, centralized, effective government, chaos would reign in the state of nature.

3. Wittgenstein’s first work was the Tractatus-Logico Philosophicus.

4. Mill’s theory of morality is known as Utilitarianism

5. According to Kant, we should always always always follow the categorical imperative.

6. How much wood would a woodchuck chuck, if a woodchuck would chuck, wood?

Section 2. True/False (5 pts each)

Ho, hum, another batch of instructions. But this time we will make them a little shorter, because we assume that students realize what they should do. \overline{False} All animals are created equal, but some animals are more equal than others. \overline{False} Laden swallows fly faster than unladen swallows, unless they carry coconuts. \overline{True} ‘Monty Python and the Holy Grail’ is a very funny movie. \overline{True} This sentence is not false. \overline{False} ‘Roger & Me’ chronicles one man’s attempt to get into Disneyland so that he can visit Toontown.

Section 3.

These are meant to be multiple-choice questions — the type you would give students to fill out using scantrons (or whatever you call the number-2 pencil automatic grading machines).

1.

How many people live in Wales?

(a) Approximately 2,811,865.

(b) More than in most countries.

(c) None.

(d) Exactly seventeen. 2.

What philosopher declared that human beings are the conduit through which language passes?

(a) Martha Stewart

(b) Jacques Derrida

(c) Jacques Cousteau

(d) Martin Heidegger 3.

The official state motto of Idaho is?

(a) “We’re friendly than other states.”

(b) “Home of the Unabomber.”

(c) “Famous Potatoes.”

(d) “Try us, you’ll like us.” 4.

How many cows does it take to graze a field?

(a) One.

(b) Two.

(c) Three.

(d) Four

Section 4. Short Answer (10 pts each)

This is an example of the shortanswer type of question, where the questions are rearranged between tests.

1. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back. Answer: Ask Lewis Carroll.

2. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy. Answer: Ask Lewis Carroll.

3. ‘Beware the Jabberwock, my son! The jaws that bite, the claws that catch! Beware the Jubjub bird, and shun the frumious Bandersnatch!’ Answer: Ask Lewis Carroll.

4. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; All mimsy were the borogoves, and the mome raths outgrabe. Answer: Ask Lewis Carroll.

5. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe. Answer: Ask Lewis Carroll.

6. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came! Answer: Ask Lewis Carroll.

7. He took his vorpal sword in hand: long time the manxome foe he sought – so rested he by the Tumtum tree, and stood awhile in thought. Answer: Ask Lewis Carroll.

Section 5. Essay Questions (50 pts each)

This is an example of the fixed type of question, where the questions aren’t rearranged between tests. Perfect for essay questions, which you normally put at the end and don’t need to rearrange.

1. Question 1.

Answer: The answer to question 1.

2. Question 2. Answer: The answer to question 2.

3. Question 3. Answer: The answer to question 3.

4. Question 4. Answer: The answer to question 4.

5. Question 5. Answer: The answer to question 5.

Philosophy 29: Critical Reasoning

Exam, Form: B Name:

Student Number:

TA:

Date:

Section 1. Fill in the blank (5 pts each)

Here is a plaçe where you can put some instructions, so that the students won’t get confused when they see a piece of paper with a whole bunch of questions on it. Of course, the instructions aren’t required to be that lengthy, but you can make them as lengthy as you want so that people know what you are talking about.

1. One main component of Nietzche’s moral philosophy is the .

2. thought that without a strong, centralized, effective government, chaos would reign in the state of nature.

3.  ’s first work was the Tractatus-  Philosophicus.

4. According to Kant, we should always always always follow the imperative.

5. Mill’s theory of morality is known as

6. How much  would a  chuck, if a  would  , wood?

Section 2. True/False (5 pts each)

Ho, hum, another batch of instructions. But this time we will make them a little shorter, because we assume that students realize what they should do.

‘Monty Python and the Holy Grail’ is a very funny movie.

This sentence is not false.

All animals are created equal, but some animals are more equal than others.

‘Roger & Me’ chronicles one man’s attempt to get into Disneyland so that he can visit Toontown.

Laden swallows fly faster than unladen swallows, unless they carry coconuts.

Section 3.

These are meant to be multiple-choice questions — the type you would give students to fill out using scantrons (or whatever you call the number-2 pencil automatic grading machines).

1.

What philosopher declared that human beings are the conduit through which language passes?

(a) Martha Stewart

(b) Jacques Derrida

(c) Jacques Cousteau

(d) Martin Heidegger

2.

The official state motto of Idaho is?

(a) “We’re friendly than other states.”

(b) “Home of the Unabomber.”

(c) “Famous Potatoes.”

(d) “Try us, you’ll like us.”

3.

How many people live in Wales?

(a) Approximately 2,811,865.

(b) More than in most countries.

(c) None.

(d) Exactly seventeen. 4.

How many cows does it take to graze a field?

(a) One.

(b) Two.

(c) Three.

(d) Four

Section 4. Short Answer (10 pts each)

This is an example of the shortanswer type of question, where the questions are rearranged between tests.

1. ‘Beware the Jabberwock, my son! The jaws that bite, the claws that catch! Beware the Jubjub bird, and shun the frumious Bandersnatch!’

2. He took his vorpal sword in hand: long time the manxome foe he sought – so rested he by the Tumtum tree, and stood awhile in thought.

3. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came!

4. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; All mimsy were the borogoves, and the mome raths outgrabe.

5. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back.

6. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe.

7. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy.

Section 5. Essay Questions (50 pts each)

This is an example of the fixed type of question, where the questions aren’t rearranged between tests. Perfect for essay questions, which you normally put at the end and don’t need to rearrange.

1. Question 1.

2. Question 2.

3. Question 3.

4. Question 4.

5. Question 5.

Answer Key for Exam B

Section 1. Fill in the blank (5 pts each)

Here is a plaçe where you can put some instructions, so that the students won’t get confused when they see a piece of paper with a whole bunch of questions on it. Of course, the instructions aren’t required to be that lengthy, but you can make them as lengthy as you want so that people know what you are talking about.

1. One main component of Nietzche’s moral philosophy is the will to power.

2. Hobbes thought that without a strong, centralized, effective government, chaos would reign in the state of nature.

3. Wittgenstein’s first work was the Tractatus-Logico Philosophicus.

4. According to Kant, we should always always always follow the categorical imperative.

5. Mill’s theory of morality is known as Utilitarianism

6. How much wood would a woodchuck chuck, if a woodchuck would chuck, wood?

Section 2. True/False (5 pts each)

Ho, hum, another batch of instructions. But this time we will make them a little shorter, because we assume that students realize what they should do.

\overline{True} ‘Monty Python and the Holy Grail’ is a very funny movie. \overline{True} This sentence is not false. \overline{False} All animals are created equal, but some animals are more equal than others. \overline{False} ‘Roger & Me’ chronicles one man’s attempt to get into Disneyland so that he can visit Toontown. \overline{False} Laden swallows fly faster than unladen swallows, unless they carry coconuts.

Section 3.

These are meant to be multiple-choice questions — the type you would give students to fill out using scantrons (or whatever you call the number-2 pencil automatic grading machines).

1.

What philosopher declared that human beings are the conduit through which language passes?

(a) Martha Stewart

(b) Jacques Derrida

(c) Jacques Cousteau

(d) Martin Heidegger 2.

The official state motto of Idaho is?

(a) “We’re friendly than other states.”

(b) “Home of the Unabomber.”

(c) “Famous Potatoes.”

(d) “Try us, you’ll like us.”

3.

How many people live in Wales?

(a) Approximately 2,811,865.

(b) More than in most countries.

(c) None.

(d) Exactly seventeen. 4.

How many cows does it take to graze a field?

(a) One.

(b) Two.

(c) Three.

(d) Four

Section 4. Short Answer (10 pts each)

This is an example of the shortanswer type of question, where the questions are rearranged between tests.

1. ‘Beware the Jabberwock, my son! The jaws that bite, the claws that catch! Beware the Jubjub bird, and shun the frumious Bandersnatch!’ Answer: Ask Lewis Carroll.

2. He took his vorpal sword in hand: long time the manxome foe he sought – so rested he by the Tumtum tree, and stood awhile in thought. Answer: Ask Lewis Carroll.

3. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came! Answer: Ask Lewis Carroll.

4. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; All mimsy were the borogoves, and the mome raths outgrabe. Answer: Ask Lewis Carroll.

5. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back. Answer: Ask Lewis Carroll.

6. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe. Answer: Ask Lewis Carroll.

7. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy. Answer: Ask Lewis Carroll.

Section 5. Essay Questions (50 pts each)

This is an example of the fixed type of question, where the questions aren’t rearranged between tests. Perfect for essay questions, which you normally put at the end and don’t need to rearrange.

1. Question 1.

Answer: The answer to question 1.

2. Question 2. Answer: The answer to question 2.

3. Question 3. Answer: The answer to question 3.

4. Question 4. Answer: The answer to question 4.

5. Question 5. Answer: The answer to question 5.

Philosophy 29: Critical Reasoning

Exam, Form: C Name:

Student Number:

TA:

Date:

Section 1. Fill in the blank (5 pts each)

Here is a plaçe where you can put some instructions, so that the students won’t get confused when they see a piece of paper with a whole bunch of questions on it. Of course, the instructions aren’t required to be that lengthy, but you can make them as lengthy as you want so that people know what you are talking about.

1. How much  would a  chuck, if a  would  , wood?

2. One main component of Nietzche’s moral philosophy is the .

3. According to Kant, we should always always always follow the imperative.

4. thought that without a strong, centralized, effective government, chaos would reign in the state of nature.

5. Mill’s theory of morality is known as

6.  ’s first work was the Tractatus-  Philosophicus.

Section 2. True/False (5 pts each)

Ho, hum, another batch of instructions. But this time we will make them a little shorter, because we assume that students realize what they should do.

‘Roger & Me’ chronicles one man’s attempt to get into Disneyland so that he can visit Toontown.

All animals are created equal, but some animals are more equal than others.

This sentence is not false.

‘Monty Python and the Holy Grail’ is a very funny movie.

Laden swallows fly faster than unladen swallows, unless they carry coconuts.

Section 3.

These are meant to be multiple-choice questions — the type you would give students to fill out using scantrons (or whatever you call the number-2 pencil automatic grading machines).

1.

How many people live in Wales?

(a) Approximately 2,811,865.

(b) More than in most countries.

(c) None.

(d) Exactly seventeen.

2.

The official state motto of Idaho is?

(a) “We’re friendly than other states.”

(b) “Home of the Unabomber.”

(c) “Famous Potatoes.”

(d) “Try us, you’ll like us.”

3.

How many cows does it take to graze a field?

(a) One.

(b) Two.

(c) Three.

(d) Four 4.

What philosopher declared that human beings are the conduit through which language passes?

(a) Martha Stewart

(b) Jacques Derrida

(c) Jacques Cousteau

(d) Martin Heidegger

Section 4. Short Answer (10 pts each)

This is an example of the shortanswer type of question, where the questions are rearranged between tests.

1. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; All mimsy were the borogoves, and the mome raths outgrabe.

2. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back.

3. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy.

4. ‘Beware the Jabberwock, my son! The jaws that bite, the claws that catch! Beware the Jubjub bird, and shun the frumious Bandersnatch!’

5. He took his vorpal sword in hand: long time the manxome foe he sought – so rested he by the Tumtum tree, and stood awhile in thought.

6. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came!

7. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe.

Section 5. Essay Questions (50 pts each)

This is an example of the fixed type of question, where the questions aren’t rearranged between tests. Perfect for essay questions, which you normally put at the end and don’t need to rearrange.

1. Question 1.

2. Question 2.

3. Question 3.

4. Question 4.

5. Question 5.

Answer Key for Exam C

Section 1. Fill in the blank (5 pts each)

Here is a plaçe where you can put some instructions, so that the students won’t get confused when they see a piece of paper with a whole bunch of questions on it. Of course, the instructions aren’t required to be that lengthy, but you can make them as lengthy as you want so that people know what you are talking about.

1. How much wood would a woodchuck chuck, if a woodchuck would chuck, wood?

2. One main component of Nietzche’s moral philosophy is the will to power.

3. According to Kant, we should always always always follow the categorical imperative.

4. Hobbes thought that without a strong, centralized, effective government, chaos would reign in the state of nature.

5. Mill’s theory of morality is known as Utilitarianism

6. Wittgenstein’s first work was the Tractatus-Logico Philosophicus.

Section 2. True/False (5 pts each)

Ho, hum, another batch of instructions. But this time we will make them a little shorter, because we assume that students realize what they should do. \overline{False} ‘Roger & Me’ chronicles one man’s attempt to get into Disneyland so that he can visit Toontown. \overline{False} All animals are created equal, but some animals are more equal than others. \overline{True} This sentence is not false. \overline{True} ‘Monty Python and the Holy Grail’ is a very funny movie. \overline{False} Laden swallows fly faster than unladen swallows, unless they carry coconuts.

Section 3.

These are meant to be multiple-choice questions — the type you would give students to fill out using scantrons (or whatever you call the number-2 pencil automatic grading machines).

1.

How many people live in Wales?

(a) Approximately 2,811,865.

(b) More than in most countries.

(c) None.

(d) Exactly seventeen. 2.

The official state motto of Idaho is?

(a) “We’re friendly than other states.”

(b) “Home of the Unabomber.”

(c) “Famous Potatoes.”

(d) “Try us, you’ll like us.”

3.

How many cows does it take to graze a field?

(a) One.

(b) Two.

(c) Three.

(d) Four 4.

What philosopher declared that human beings are the conduit through which language passes?

(a) Martha Stewart

(b) Jacques Derrida

(c) Jacques Cousteau

(d) Martin Heidegger

Section 4. Short Answer (10 pts each)

This is an example of the shortanswer type of question, where the questions are rearranged between tests.

1. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; All mimsy were the borogoves, and the mome raths outgrabe. Answer: Ask Lewis Carroll.

2. One, two! One, two! And through and through the vorpal blade went snicker-snack! He left it dead, and with its head he went galumphing back. Answer: Ask Lewis Carroll.

3. ‘And has thou slain the Jabberwock? Come to my arms, my beamish boy! O frabjous day! Calooh! Callay!’ He chortled in his joy. Answer: Ask Lewis Carroll.

4. ‘Beware the Jabberwock, my son! The jaws that bite, the claws that catch! Beware the Jubjub bird, and shun the frumious Bandersnatch!’ Answer: Ask Lewis Carroll.

5. He took his vorpal sword in hand: long time the manxome foe he sought – so rested he by the Tumtum tree, and stood awhile in thought. Answer: Ask Lewis Carroll.

6. And as in uffish thought he stood, the Jabberwock, with eyes of flame, came whiffling through the tulgey wood, and burbled as it came! Answer: Ask Lewis Carroll.

7. ‘Twas brillig, and the slithy toves did gyre and gimble in the wabe; all mimsy were the borogoves, and the mome raths outgrabe. Answer: Ask Lewis Carroll.

Section 5. Essay Questions (50 pts each)

This is an example of the fixed type of question, where the questions aren’t rearranged between tests. Perfect for essay questions, which you normally put at the end and don’t need to rearrange.

1. Question 1.

Answer: The answer to question 1.

2. Question 2. Answer: The answer to question 2.

3. Question 3. Answer: The answer to question 3.

4. Question 4. Answer: The answer to question 4.

5. Question 5. Answer: The answer to question 5.